# Assessing Language Comprehension in Large Language Models Using Construction Grammar

### v1

@@ Appendix A Construction Grammers
CxG has particular explanatory power with respect to phrasal constructions, such as the resultative construction: “The jackhammer pounded us deaf.” Generative linguistic theories (e.g., Chomsky (2014)) would generally analyze a transitive sentence with one verbal head (“pounded”) that licenses the arguments of the sentence. “Pounded” generally licenses an agent subject (here, “jackhammer”) and potentially a patient direct object. Generative approaches argue that this information about the verb is memorized and stored in the lexicon, while combinatory rules of how to put lexical items together are stored in a separate syntax module of language processing. However, unless a special grammatical rule or sense of the verb is postulated, there is nothing to explain why “us” is not the thing pounded here, or what licenses the adjective “deaf.” Nonetheless, native speakers have no problem recognizing the special formal and semantic properties of this Cxn, namely that it entails a pounding event that causes a change in state of “us” resulting in the state of “deaf.”
@@
In contrast to Generative linguistic theories (e.g., Chomsky (2014)), CxG posits that speakers acquire and store constructions, which notably account for not only the semantic properties of the unit but also the formal syntactic properties. Constructions at all levels of language are learned through language usage; thus, in lieu of grammatical rules accounting for the grammaticality of a particular structure, frequency plays an important role in what ‘sounds right’ or is grammatical to a speaker. In the CxG account, speakers acquire single-word constructions that they are frequently exposed to (e.g., “milk”), but then generalize from that to recognize how an acquired holophrastic Cxn falls into slots of larger, more complex constructions (e.g., “want milk,” “baby want milk,” “want up,” “I want to go”). In this process of generalization, speakers build up a taxonomically organized set of the related constructions of their language, or a constructicon. Speakers use domain-general cognitive processes to incrementally generalize over such frequent constructions to novel usages, arriving at the ability to interpret even rare and previously unseen instances of a Cxn.
Appendix B Templatic Constructional NLI vs Freeform NLI
| Setting | IC Data | Dataset | Accuracy | |
| GPT-3.5 | GPT-4o | |||
| Zero-shot | None | Templatic | 0.6 | 0.88 |
| One-shot | CxNLI | Templatic | 0.69 | 0.9 |
| Three-shot | CxNLI | Templatic | 0.79 | 0.96 |
| One-shot | SNLI | Templatic | 0.58 | 0.86 |
| Three-shot | SNLI | Templatic | 0.59 | 0.86 |
[…]
| One-shot | CxNLI | Freeform | 0.78 | 0.91 |
| Three-shot | CxNLI | Freeform | 0.83 | 0.94 |
| One-shot | SNLI | Freeform | 0.7 | 0.89 |
| Three-shot | SNLI | Freeform | 0.69 | 0.9 |
@@ Appendix B Templatic Constructional NLI vs Freeform NLI
In addition to our templatically generated constructional NLI dataset, we construct a more "freeform" NLI dataset by training a person to create hypotheses that are intended to be more similar in form to hypotheses from SNLI. Both our "freeform" and templatic datasets are constructed with real-world corpus data of constructions, primarily coming from the COGS dataset (Bonial and Tayyar Madabushi, 2024), with supplementary data coming from Corpus of Contemporary American English (COCA, Davies 2010), and the English Corpus from the Web, or EnCOW (Schäfer and Bildhauer, 2012; Schäfer, 2015). Within both of these datasets, each premise includes one of 8 total constructions: the comparative-correlative Cxn, the let-alone construction, the way-manner construction, the causative-with construction, the conative construction, the resultative construction, the caused motion construction, and the intransitive motion Cxn. We include a roughly balanced sample of each construction, with all premises taken from corpus data. These constructions cover a wide range of schematicity meaning that they have different levels of lexicalization/abstractness.
@@
| Dataset | Ex. | Tokens | IAA |
| Cxn NLI (Freeform) | 435 | 15,401 | 92% |
| Cxn NLI (Templatic) | 435 | 15,144 | 90% |
| challenge NLI | 99 | 3,202 | 82% |
| challenge Reasoning | 740 | 41,835 | 78% |
| Prompt Type | Accuracy | |
| GPT-3.5 | GPT-4o | |
| Variation 1 | 0.74 | 0.95 |
| Variation 2 | 0.79 | 0.92 |
| Variation 3 | 0.76 | 0.94 |
| Variation 4 | 0.74 | 0.93 |
[…]
| Best Variation Rephrase 2 | 0.67 | 0.95 |
| Best Variation Rephrase 3 | 0.69 | 0.96 |
| Prompt Type | Accuracy | |
| GPT-3.5 | GPT-4o | |
| Variation 1 | 0.24 | 0.53 |
| Variation 2 | 0.28 | 0.46 |
| Variation 3 | 0.31 | 0.57 |
| Variation 4 | 0.26 | 0.54 |
| Best Variation Rephrase 1 | 0.31 | 0.57 |
| Best Variation Rephrase 2 | 0.26 | 0.55 |
| Best Variation Rephrase 3 | 0.23 | 0.54 |
| Prompt Type | Accuracy | |
| GPT-3.5 | GPT-4o | |
| Variation 1 | 0.41 | 0.61 |
| Variation 2 | 0.38 | 0.61 |
| Variation 3 | 0.45 | 0.64 |
| Variation 4 | 0.42 | 0.68 |
[…]
| Best Variation Rephrase 2 | 0.58 | 0.68 |
| Best Variation Rephrase 3 | 0.52 | 0.68 |
| Cxn Name | Cxn NLI (Freeform) | Cxn NLI (Templatic) | |
@@
| Causative-With | Premise | Freshly ground coffee beans filled the room with a seductive, earthy aroma. | Freshly ground coffee beans filled the room with a seductive, earthy aroma. |
| Hypothesis | Stale whole coffee beans filled the room with an unattractive, otherworldly scent. | The room did not contain a seductive, earthy aroma. | |
| Relation | Contradiction | Contradiction | |
| Caused-Motion | Premise | I threw the stone across the river. | I threw the stone across the river. |
| Hypothesis | I threw the stone up and caught it again. | I caused the stone to move across the river by throwing it. | |
| Relation | Contradiction | Entailment | |
@@
| Comparative-Correlative | Premise | The more they work, the more I will pay them. | The more they work, the more I will pay them. |
@@
| Hypothesis | I intend to compensate my workers based on their amount of work. | Increasing the amount they work will increase the amount I pay them. | |
| Relation | Entailment | Entailment | |
| Conative | Premise | I sipped at the Heineken. | I sipped at the Heineken. |
| Hypothesis | I spit out the Heineken. | The Heineken was not the target of my sipping. | |
| Relation | Contradiction | Contradiction | |
| Intransitive-Motion | Premise | I ran around the track. | I ran around the track. |
| Hypothesis | I walked around the track. | I ran, staying in one place. | |
| Relation | Contradiction | Contradiction | |
@@
| Let-Alone | Premise | It’s unsurprising that such an attitude failed to produce competent screenwriters, let alone exciting ones. | It’s unsurprising that such an attitude failed to produce competent screenwriters, let alone exciting ones. |
@@
| Hypothesis | Although such an attitude failed to produce competent screenwriters, it produced plenty of exciting screenwriters. | An attitude that produces exciting screenwriters can also produce competent ones. | |
| Relation | Contradiction | Entailment | |
| Resultative | Premise | The jackhammer pounded us deaf. | The jackhammer pounded us deaf. |
| Hypothesis | The jackhammer was easy on our ears. | We were completely deaf before the jackhammer pounded. | |
| Relation | Contradiction | Contradiction | |
| Way-Manner | Premise | I yawned my way back to the Narrow Neck. | I yawned my way back to the Narrow Neck. |
| Hypothesis | I yawned at least once on my journey back to the Narrow Neck. | I traveled back to Narrow Neck without yawning. | |
| Relation | Entailment | Contradiction |
The freeform Constructional NLI tuples are constructed in order to be similar to other NLI datasets, testing general types of semantics. These tuples do not always necessitate perfect understanding of the present Cxn in order to be correct; rather, the surrounding context is sometimes enough to provide the correct answer. On the other hand, the templatic Constructional NLI dataset is designed to specifically test the meanings associated with the present constructions. Each NLI tuple makes direct use of the meaning attached with the constructions, and thus providing the correct answer to the NLI task implies correctly associating some aspect of meaning with the Cxn. These templates also allow us to easily construct our challenge examples, which is crucial for the experiment laid out in 5. Examples of Cxn and NLI templates are provided in Table 7. For full examples of NLI triples for each of the 8 constructions evaluated, see Table 12. We test GPT-4o and GPT-3.5 on our freeform Cxn NLI dataset, and find generally similar trends to the templatic Cxn NLI dataset. A results comparison across the two NLI datasets is shown in Table 6.
Appendix C Scientific Artifacts and Descriptive Statistics for All Datasets
We use the following scientific artifacts: COCA (Davies, 2010), EnCOW (Schäfer and Bildhauer, 2012; Schäfer, 2015), the CoGS dataset (Bonial and Tayyar Madabushi, 2024), and the OpenAI API, in addition to our created datasets. The COCA corpus contains 8 genres: Academic, Blog, Fiction, Magazine, News, Spoken, TV, and Web. It is intended to capture American English, though there is no guarantee that it does not also include some other varieties. Demographic information about the creators of the texts in the corpus is not always available given its scale. EnCOW is a large scale corpus of English text from the web. As such, the demographic information that it captures is not completely clear. All datasets besides COCA and EnCOW are open-source under a Creative Commons license. The institutions of the authors have valid licenses for COCA and EnCOW, permitting their use in academic settings. We use the artifacts as intended. Overall, we construct and experiment on 4 datasets: 2 constructional NLI datasets (templatic and freeform), the challenge NLI dataset, and the challenge reasoning dataset. The sizes and IAA agreement for the final datasets is reported in Table 8. All of our datasets are exclusively in English.
Appendix D Prompt Variation Experiments
@@
For each experiment, we choose one setting to observe the impact of prompt variations on performance.121212For NLI we choose the 3-shot CxNLI setting. For challenge reasoning we choose zero-shot with a subset of all constructions. These prompt variations were explored in two sequential stages; changing the content of the prompt and changing the phrasing of the prompt. Changing the content of the prompt involved giving more details about the task, giving fewer details about the task and experimenting with the specific instructions. After selecting the best-performing prompt content, the prompt was rephrased using an LLM and variations of the prompt with new wording were used to check the robustness of the model output by observing how small variations in the prompts affect performance. We report an example of each of our prompts for variations in content in Tables 13 and 14. Results for our prompt variation experiments are shared in Tables 9, 10 and 11. Annotators were shown the exact same prompts as models when doing their annotations.
@@
| Prompt Type | Prompt |
| Variation 1 | You are the world’s best annotator. Your task is to read questions from a dataset, and determine the correct answer from the options provided using only the information given in the question. Output a single character of either A, B or C and nothing else. |
| Variation 2 | Using only the information given in the question, output a single character of either A, B, or C, corresponding to the correct answer. Output a single character only and nothing else. |
| Variation 3 | You are the world’s best annotator. Your task is to read questions from a dataset, and determine the correct answer from the options provided using only the information given in the question. Output a single character of either A, B or C and nothing else. If the answer is not clear from the context provided, output the character C. |
| Variation 4 | Using only the information given in the question, output a single character of either A, B, or C, corresponding to the correct answer. Output a single character only and nothing else. If the answer is not clear from the context provided, output the character C. |
Appendix E Example NLI Tuples
In Table 12, we show examples of our constructional NLI datasets. We show challenge Cxn NLI examples in Table 15.
| CxN Name | Example NLI Triple | |
| Conative | Premise | I watch the women, their legs crossed at the ankles, try to look as if they don’t sweat at all. |
| Hypothesis | Their legs made a crossing motion towards their ankles. | |
| Relation | Contradiction | |
| Causative- with | Hypothesis | He hit the lamp with his head. |
| Premise | He caused the lamp to contain his head. | |
| Relation | Contradiction | |
| Intransitive- motion | Hypothesis | He spoke to the workers on the street corner. |
| Premise | He changed locations by speaking. | |
| Relation | Contradiction | |
| Caused- motion | Hypothesis | I introduced her to my boss. |
| Premise | She remained in the same place. | |
| Relation | Neutral | |
| Resultative | Hypothesis | A famous emperor buried China’s scholars alive with their books. |
| Premise | Burying caused the scholars to become alive. | |
| Relation | Contradiction |
@@ Appendix F Chain-of-Thought Results for all Experiments
We replicate each of our main experiments with the addition of Chain-of-Thought (CoT) to observe changes in performance and inspect any errors in the model’s reasoning steps. For CoT thought prompting we do not provide examples of reasoning in the prompt, instead we simply ask the model to explain "step by step". The results for each experiment have been shared in tables 16, 17, 18 and 19.
@@
| Zero-shot | None | 0.62 | 0.89 |
| One-shot | Challenge | 0.6 | 0.9 |
| One-shot | Challenge, CxNLI | 0.6 | 0.9 |
| Setting | IC Data | Accuracy | |
| GPT-3.5 | GPT-4o | ||
@@
| Cxn | Accuracy | |
| GPT-3.5 | GPT-4o | |
| comparative correlative | 0.51 | 0.68 |
| let-alone | 0.51 | 0.63 |
| way manner | 0.63 | 0.64 |
| all | 0.53 | 0.66 |
Appendix G Reasoning Test Description and Examples
There are 7 bias and calibration tests outlined in Weissweiler et al. (2022) for the comparative-correlative. However, their tests are templatically generated using one specific syntactic form of the comparative-correlative. We adapt these tests to natural examples of the construction, and apply the tests to let-alone and way-manner. We also introduce a new test called "Nonce Calibration", which is similar in spirit to the "Adjective Calibration" test but using a nonce word. A brief description of all test types is below.
1.
Base - this test presents an example of a construction, along with additional context, followed by a question. The constructional meaning combines with the context to provide the answer to the question.
2.
Recency Bias - this test reorders the Cxn and the contextual information, placing the contextual information first. The answer should be the same as Base.
3.
Vocabulary Bias - this test reverses the semantics of the scalar constructions by swapping vocabulary items in different slots of the Cxn. The answer should be the opposite of Base.131313This test is possible for ”Let-alone” and ”Comparative-Correlative”, but is excluded for ”Way-Manner”
4.
Name Bias - this test simply changes the nouns that are referenced in the context of the construction, to see if the model is biased towards interpreting certain Cxns differently depending on the nouns that fill their slots. The answer should be the same as Base.
5.
Short Calibration - this test removes the Cxn altogether, making it impossible to answer the question. The model should respond that it does not have enough information to answer.
6.
Name Calibration - this test gives contextual information unrelated to the construction, in the form of nouns unrelated to the question. The model should respond that it does not have enough information to answer.
7.
Adjective Calibration - this test again provides irrelevant contextual information, in the form of an irrelevant adjective. The model should respond that it does not have enough information to answer.
8.
Nonce Calibration - this test provides irrelevant contextual information in the form of a nonce word instead of meaningful contextual information. This test is the clearest example of a case where the model has no prior reason to provide any answer, and the model should respond that it does not have enough information.
Examples for each of the reasoning tests for each of our constructions can be found below in Table 20.
| Cxn Name | Test Type | Test Text | Answer |
| Comparative Correlative | Base | The more they work, the more I will pay them. The less they work, the less I will pay them. They work more in December than they do in September. How much will I pay them in December, compared to in September? | pay more |
[…]
| Let Alone | Base | Beecher’s reputation as a preacher, let alone as a Man of God, was not universally accepted. Beecher’s reputation as a preacher was weaker in July than in December. When was Beecher more likely to be seen as a Man of God, December or July? | December |
[…]
| Way-Manner | Base | Thus a miner who slipped from the worn and greasy rungs was liable to fall a long way and a dropped tool could clatter its way down to cause death or injury to some fellow below. A dropped tool did more clattering on Monday than on Saturday. When was the dropped tool likely to be falling, on Monday or on Saturday? | Monday |
[…]
Appendix H Annotator Information
All datasets are annotated by co-authors of this paper. 2 annotators identify as men, 4 annotators identify as women. Of our annotators, 3 have a graduate degree in linguistics while 3 do not. At least one linguistic expert and one non-expert annotate each dataset. Because the source of our datasets is web corpora, there is some risk of offensive or hateful content. During annotation, annotators were asked to remove any hateful or offensive content as well as personal identifying information. The annotation task was given to annotators in a spreadsheet. The instructions were identical to model prompts, to ensure maximum comparability to models. All annotators were generally familiar with the tasks of NLI and QA in general, so no general task-related training was needed.
Appendix I Model Hyperparameters

### v2

@@ Appendix A Construction Grammars
CxG has particular explanatory power with respect to phrasal Cxns, such as the resultative Cxn: “The jackhammer pounded us deaf.” Generative linguistic theories (e.g., Chomsky (2014b)) would generally analyze a transitive sentence with one verbal head (“pounded”) that licenses the arguments of the sentence. “Pounded” generally licenses an agent subject (here, “jackhammer”) and potentially a patient direct object. Generative approaches argue that this information about the verb is memorized and stored in the lexicon, while combinatory rules of how to put lexical items together are stored in a separate syntax module of language processing. However, unless a special grammatical rule or sense of the verb is postulated, there is nothing to explain why “us” is not the thing pounded here, or what licenses the adjective “deaf.” Nonetheless, native speakers have no problem recognizing the special formal and semantic properties of this Cxn, namely that it entails a pounding event that causes a change in state of “us” resulting in the state of “deaf.”
@@
In contrast to Generative linguistic theories (e.g., Chomsky (2014b)), CxG posits that speakers acquire and store Cxns, which notably account for not only the semantic properties of the unit but also the formal syntactic properties. Cxns at all levels of language are learned through language usage; thus, in lieu of grammatical rules accounting for the grammaticality of a particular structure, frequency plays an important role in what ‘sounds right’ or is grammatical to a speaker. In the CxG account, speakers acquire single-word Cxns that they are frequently exposed to (e.g., “milk”), but then generalize from that to recognize how an acquired holophrastic Cxn falls into slots of larger, more complex Cxns (e.g., “want milk,” “baby want milk,” “want up,” “I want to go”). In this process of generalization, speakers build up a taxonomically organized set of the related Cxns of their language, or a constructicon. Speakers use domain-general cognitive processes to incrementally generalize over such frequent Cxns to novel usages, arriving at the ability to interpret even rare and previously unseen instances of a Cxn.
Appendix B Scientific Artifacts and Descriptive Statistics for All Datasets
We use the following scientific artifacts: COCA (Davies, 2010), EnCOW (Schäfer and Bildhauer, 2012; Schäfer, 2015), the CoGS dataset (Bonial and Tayyar Madabushi, 2024b), and the OpenAI API, in addition to our created datasets. The COCA corpus contains 8 genres: Academic, Blog, Fiction, Magazine, News, Spoken, TV, and Web. It is intended to capture American English, though there is no guarantee that it does not also include some other varieties. Demographic information about the creators of the texts in the corpus is not always available given its scale. EnCOW is a large scale corpus of English text from the web. As such, the demographic information that it captures is not completely clear. All datasets besides COCA and EnCOW are open-source under a Creative Commons license. The institutions of the authors have valid licenses for COCA and EnCOW, permitting their use in academic settings. We use the artifacts as intended. Overall, we construct and experiment on 2 datasets: CxNLI (Exp1) and CxNLI-Distinction (Exp2). The sizes and IAA agreement for the final datasets is reported in Table 7. All of our datasets are exclusively in English.
| Dataset | N | Tokens | IAA |
| CxNLI (Exp1) | 435 | 15,144 | 90% |
| CxNLI-Distinction (Exp2) | 99 | 3,202 | 83% |
B.1 CxNLI (Exp 1) Data Sources
@@
The templatic datasets is constructed with real-world corpus data of Cxns, primarily coming from the COGS dataset (Bonial and Tayyar Madabushi, 2024b), with supplementary data coming from Corpus of Contemporary American English (COCA, Davies 2010), and the English Corpus from the Web, or EnCOW (Schäfer and Bildhauer, 2012; Schäfer, 2015). Within this dataset, each premise includes one of 8 total Cxns: the comparative-correlative Cxn, the let-alone Cxn, the way-manner Cxn, the causative-with Cxn, the conative Cxn, the resultative Cxn, the caused motion Cxn, and the intransitive motion Cxn. We include a roughly balanced sample of each Cxn, with all premises taken from corpus data. These Cxns cover a wide range of schematicity meaning that they have different levels of lexicalization/abstractness.
Appendix C Annotator Information
All datasets are annotated by co-authors of this paper. 2 annotators identify as men, 4 annotators identify as women. Of our annotators, 3 have a graduate degree in linguistics while 3 do not. At least one linguistic expert and one non-expert annotate each dataset. Because the source of our datasets is web corpora, there is some risk of offensive or hateful content. During annotation, annotators were asked to remove any hateful or offensive content as well as personal identifying information. The annotation task was given to annotators in a spreadsheet. The instructions provided are detailed in Appendix H.
Appendix D Example NLI Tuples
Our templates for generating hypotheses across Cxn types is shown in Table 8. In Table 9, we show examples of our constructional NLI datasets from Exp 1. We show examples of the distinction-requiring NLI examples from Exp 2, alongside examples of our new constructions for Exp 2, in Table 10.
@@ Appendix D Example NLI Tuples
| Cxn Name | CxNLI (Exp 1) | |
@@
| Causative-With | Premise | Freshly ground coffee beans filled the room with a seductive, earthy aroma. |
| Hypothesis | The room did not contain a seductive, earthy aroma. | |
| Relation | Contradiction | |
| Caused-Motion | Premise | I threw the stone across the river. |
| Hypothesis | I caused the stone to move across the river by throwing it. | |
| Relation | Entailment | |
@@
| Comparative-Correlative | Premise | The more they work, the more I will pay them. |
@@
| Hypothesis | Increasing the amount they work will increase the amount I pay them. | |
| Relation | Entailment | |
| Conative | Premise | I sipped at the Heineken. |
| Hypothesis | The Heineken was not the target of my sipping. | |
| Relation | Contradiction | |
| Intransitive-Motion | Premise | I ran around the track. |
| Hypothesis | I ran, staying in one place. | |
| Relation | Contradiction | |
@@
| Let-Alone | Premise | It’s unsurprising that such an attitude failed to produce competent screenwriters, let alone exciting ones. |
@@
| Hypothesis | An attitude that produces exciting screenwriters can also produce competent ones. | |
| Relation | Entailment | |
| Resultative | Premise | The jackhammer pounded us deaf. |
| Hypothesis | We were completely deaf before the jackhammer pounded. | |
| Relation | Contradiction | |
| Way-Manner | Premise | I yawned my way back to the Narrow Neck. |
| Hypothesis | I traveled back to Narrow Neck without yawning. | |
| Relation | Contradiction |
| CxN Name | Example NLI Triple | ||
| Intransitive+At | Premise | I watch the women, their legs crossed at the ankles, try to look as if they don’t sweat at all. | |
| Hypothesis | Their legs made a crossing motion towards their ankles. | ||
| Relation | Contradiction | ||
Hypothesis | He hit the lamp with his head. | |
| Premise | He caused the lamp to contain his head. | ||
| Relation | Contradiction | ||
Hypothesis | He spoke to the workers on the street corner. | |
| Premise | He changed locations by speaking. | ||
| Relation | Contradiction | ||
Hypothesis | I introduced her to my boss. | |
| Premise | She remained in the same place. | ||
| Relation | Neutral | ||
| Depictive | Hypothesis | A famous emperor buried China’s scholars alive with their books. | |
| Premise | Burying caused the scholars to become alive. | ||
| Relation | Contradiction |
Appendix E Prompt Variation Experiments
@@
For each experiment, we choose one setting to observe the impact of prompt variations on performance.888For NLI we choose the 3-shot CxNLI setting. These prompt variations were explored in two sequential stages; changing the content of the prompt and changing the phrasing of the prompt. Changing the content of the prompt involved giving more details about the task, giving fewer details about the task and experimenting with the specific instructions. After selecting the best-performing prompt content, the prompt was rephrased using an LLM and variations of the prompt with new wording were used to check the robustness of the model output by observing how small variations in the prompts affect performance. We report an example of each of our prompts for variations in content in Table 11. Results for our prompt variation experiments are shared in Tables 12 and 13.
@@
| Prompt Type | Accuracy | |
| GPT-3.5 | GPT-4o | |
| Variation 1 | 0.74 | 0.95 |
| Variation 2 | 0.79 | 0.92 |
| Variation 3 | 0.76 | 0.94 |
| Variation 4 | 0.74 | 0.93 |
| Best Variation Rephrase 1 | 0.79 | 0.95 |
| Best Variation Rephrase 2 | 0.67 | 0.95 |
| Best Variation Rephrase 3 | 0.69 | 0.96 |
| Prompt Type | Accuracy | |
| GPT-3.5 | GPT-4o | |
| Variation 1 | 0.24 | 0.53 |
| Variation 2 | 0.28 | 0.46 |
| Variation 3 | 0.31 | 0.57 |
| Variation 4 | 0.26 | 0.54 |
| Best Variation Rephrase 1 | 0.31 | 0.57 |
| Best Variation Rephrase 2 | 0.26 | 0.55 |
| Best Variation Rephrase 3 | 0.23 | 0.54 |
@@ Appendix F Chain-of-Thought Results for all Experiments
We replicate each of our main experiments with the addition of Chain-of-Thought (CoT) to observe changes in performance and inspect any errors in the model’s reasoning steps. For CoT thought prompting we do not provide examples of reasoning in the prompt, instead we simply ask the model to explain "step by step". The results for each experiment have been shared in tables 14 and 15.
@@
Appendix G Model Parameters and Hyperparameters